# Circle bundles with PSC over some four manifolds

### v2

@@ 2.3. Donaldson divisors of $M\times S^{2}$
Take any regular value $p\in M$. Then the oriented count #$\#f^{-1}(p)$ is equal to the oriented intersection number #$\#([Z]\cap[\{p\}\times S^{2}])$. Therefore, we have
$\deg(f)$ =#$=\#([Z]\cap[\{p\}\times S^{2}])$
@@
The last inequality holds because $\{p\}\times S^{2}$ is a symplectic submanifold of $(M\times S^{2},\omega)$. ∎
@@ 2.4. Symplectically aspherical manifolds
Remark.
Note that it is immediate from Definition 3 that symplectic submanifolds of symplectically aspherical manifolds are also symplectically aspherical. Since symplectic manifolds $(X,\omega)$ with $\pi_{2}(X)=0$ are automatically symplectically aspherical, together with the previous observation as well as Donaldson’s theorem (Theorem 1), this provides a large class of symplectically aspherical manifolds.
@@
3. Preliminary results
In this section we assemble all other results that we will require to prove Theorem A.
3.1. Nullcobordism and macroscopic dimension
We first prove a general lemma relating macroscopic dimension of a manifold with the homotopy type of a nullcobordism of the manifold. This will be used to prove of Part (2) of Theorem A.
@@ 3.1. Nullcobordism and macroscopic dimension
To this end, let us assume without loss of generality that $f$ is a simplicial map. This can be arranged by the simplicial approximation theorem. Let $\widetilde{p}\in\widetilde{K}$ be a point lying over $p\in K$. Then, we claim $\widetilde{f}^{-1}(\widetilde{p})$ is isometric to $f^{-1}(p)$. Indeed, note that $f^{-1}(p)\hookrightarrow M$ induces the zero map on $\pi_{1}$, since otherwise there would be a loop $\gamma\subset f^{-1}(p)$ which is not nullhomotopic in $M$ such that $f_{*}[\gamma]=0\in\pi_{1}(K)$. This would be absurd, as $f$ induces an isomorphism on $\pi_{1}$. Thus, $f^{-1}(p)$ lifts homeomorphically in the cover to a unique subset in $\widetilde{Y}$ containing $\widetilde{p}$. By uniqueness, this lift is exactly $\widetilde{f}^{-1}(\widetilde{p})$. Consequently, the covering projection must be a homeomorphism:
@@
3.2. Positive scalar curvature and Donaldson divisors
We begin with the following result due to Cecchini, Räde and Zeidler in [CRZ23], which establishes a codimension-$2$ obstruction to existence of a PSC metric on a $7$-manifold.
Theorem 2.
[CRZ23, Theorem 1.11] Let $Y^{7}$ be a closed connected manifold, and $X^{5}\subset Y^{7}$ be a submanifold such that
@@ 3.2. Positive scalar curvature and Donaldson divisors
    $i:X\hookrightarrow Y$ has trivial normal bundle,
@@
    $i_{*}:\pi_{1}(X)\to\pi_{1}(Y)$ is injective,
  3. (3)
    $i_{*}:\pi_{2}(X)\to\pi_{2}(Y)$ is surjective
Then, if $X$ does not admit a PSC metric, $Y$ does not admit a PSC metric either.
Let $(M^{6},\omega)$ be an integral symplectic $6$-manifold. Let $Z\subset(M^{6},\omega)$ be a Donaldson divisor. Recall from Lemma 2.1 that the normal bundle of $Z$ in $M$ is isomorphic to $L^{\otimes k}$, which is a nontrivial complex line bundle over $Z$. However, we will now consider the unit circle bundle of $L^{\otimes k}$, denoted by $\mathbb{S}(L^{\otimes k})$. In conjunction with it, we will consider the restriction of the unit circle bundle of $L^{\otimes k}$ to $Z$, denoted by $\mathbb{S}(L^{\otimes k}|_{Z})$. Note that $\mathbb{S}(L^{\otimes k})$ is a $7$-manifold and $\mathbb{S}(L^{\otimes k}|_{Z})$ is a $5$-manifold.
In the following proposition, we shall verify that the pair $(\mathbb{S}(L^{\otimes k}),\mathbb{S}(L^{\otimes k}|_{Z}))$ satisfies the hypotheses of Theorem 2.
Proposition 7.
Let us denote $Y:=\mathbb{S}(L^{\otimes k})$ and $X:=\mathbb{S}(L^{\otimes k}|_{Z})$. Then the pair $(Y^{7},X^{5})$ satisfies the hypotheses of Theorem 2. Therefore, if $\mathbb{S}(L^{\otimes k})$ is PSC, then so is $\mathbb{S}(L^{\otimes k}|_{Z})$.
Before the proof of Proposition 7, we state and prove an elementary topology lemma.
Lemma 3.2.
Let $\pi:E\to B$ be a complex line bundle, and $\mathbb{S}(E)\subset E$ be the unit circle bundle with respect to some choice of a fiberwise Riemannian metric on $E$. Let $\pi_{s}:\mathbb{S}(E)\to B$ denote the bundle projection. Then, $\pi_{s}^{*}E$ is trivial over $\mathbb{S}(E)$.
Proof.
By definition of the pullback construction,
$\pi_{s}^{*}E=\mathbb{S}(E)\times_{B}E$
The diagonal section $\sigma:\mathbb{S}(E)\to\mathbb{S}(E)\times_{B}E$ defined by $\sigma(e):=(e,e)$ is a nowhere zero section of the circle bundle $\pi_{s}^{*}E\to\mathbb{S}(E)$. Therefore, $\pi_{s}^{*}E$ is trivial over $\mathbb{S}(E)$. ∎
Proof of Proposition 7.
First, we observe that the normal bundle of $X\subset Y$ is trivial. Indeed, let $E:=L^{\otimes k}|_{Z}$ and $\pi_{s}:\mathbb{S}(E)=X\to Z$ denote the bundle projection. Then,
$N_{Y}X=\pi^{*}N_{M}Z\cong\pi_{s}^{*}(E)$
By Lemma 3.2, we conclude $N_{Y}X$ is trivial over $\mathbb{S}(E)=X$, as desired.
Next, we show $i$ induces an isomorphism on $\pi_{1}$ and a surjection on $\pi_{2}$. To this end, consider the following commutative diagram:
[…]
The rows are fiber bundles, therefore by applying the long exact sequence in homotopy and using naturality, we have the following diagram:
00{0}${{\pi_{2}(Y)}}$${{\pi_{2}(M)}}$${{\pi_{1}(S^{1})\cong\mathbf{Z}}}$${{\pi_{1}(Y)}}$${{\pi_{1}(M)}}$00{0}00{0}${{\pi_{2}(X)}}$${{\pi_{2}(Z)}}$${{\pi_{1}(S^{1})\cong\mathbf{Z}}}$${{\pi_{1}(X)}}$${{\pi_{1}(Z)}}$00{0}$\scriptstyle{\phi_{M}}$$\scriptstyle{i_{*}}$$\scriptstyle{j_{*}}$$\scriptstyle{\phi_{Z}}$$\scriptstyle{\mathrm{id}}$$\scriptstyle{i_{*}}$$\scriptstyle{j_{*}}$
Note, as indicated in the diagram above, $j_{*}:\pi_{1}(Z)\to\pi_{1}(M)$ is an isomorphism and $j_{*}:\pi_{2}(Z)\to\pi_{2}(M)$ is a surjection due to Proposition 2 specialized to the case of $n=3$, as $M$ is a symplectic manifold of dimension $62n=$ 6.
  1. To show $i_{*}:\pi_{2}(X)\to\pi_{2}(Y)$ is surjective:
    Let $\alpha\in\pi_{2}(Y)$. Let $\alpha^{\prime}\in\pi_{2}(M)$ be the image of $\alpha$. Then $\phi_{M}(\alpha^{\prime})=0$. Since $j_{*}:\pi_{2}(Z)\to\pi_{2}(M)$ is surjective, there exists $\beta^{\prime}\in\pi_{2}(M)$ such that $j_{*}(\beta^{\prime})=\alpha^{\prime}$. By commutativity,
    $\phi_{Z}(\beta^{\prime})=\phi_{M}(j_{*}(\beta^{\prime}))=\phi_{M}(\alpha^{\prime})=0.$
    Therefore, $\beta^{\prime}$ lies in the kernel of $\phi_{Z}$, which is equal to the image of $\pi_{2}(X)$ in $\pi_{2}(Z)$. Thus, there exists $\beta\in\pi_{2}(X)$ such that image of $\beta$ is exactly $\beta^{\prime}$. Now, the image of $i_{*}(\beta)-\alpha$ in $\pi_{2}(M)$ is $j_{*}(\beta^{\prime})-\alpha^{\prime}=0$. Since $\pi_{2}(Y)$ injects into $\pi_{2}(M)$, we conclude $i_{*}(\beta)=\alpha$. This proves the claim.
  2. To show $i_{*}:\pi_{1}(X)\to\pi_{1}(Y)$ is injective:
    Note that $\phi_{M}$ and $\phi_{Z}$ have the same image and therefore the same cokernel. By the short five-lemma, $i_{*}:\pi_{1}(X)\to\pi_{1}(Y)$ is an isomorphism. In particular, it is injective. ∎
4. Proofs of the main results
4.1. Proof of Theorem A
In this section, we prove our main result. Let $(M,\eta)$ be an integral symplectic $4$-manifold. Let $\omega_{0}$ be the standard area form on $S^{2}$. We equip $M\times S^{2}$ with the integral symplectic form $\omega=\eta\oplus\omega_{0}$. Let $L$ be the pre-quantum line bundle on the symplectic manifold $(M\times S^{2},\omega)$. Let $Z\subset(M\times S^{2},\omega)$ be a Donaldson divisor, with $[Z]=\mathrm{PD}(k\cdot[\omega])$, for $k\gg 1$ sufficiently large. We set $E_{Z}:=\mathbb{S}(L^{\otimes k}|_{Z})$ as the unit circle bundle of $L^{\otimes k}$ restricted to $Z$ (as introduced in Lemma 2.1 and Proposition 7).
  1. Proof of (1).
    As $M\times S^{2}$ admits a PSC metric, and $\mathbb{S}(L^{\otimes k})$ is a circle bundle on $M\times S^{2}$, $\mathbb{S}(L^{\otimes k})$ also admits a PSC metric [BB83, Theorem C]. By Proposition 7, this implies that $E_{Z}=\mathbb{S}(L^{\otimes k}|_{Z})$ also admits a PSC metric.
  2. Proof of (2).
    Let $\nu(Z)$ denote a tubular neighbourhood of $Z\subset M\times S^{2}$. Consider the manifold with boundary $W:=(M\times S^{2})\setminus\nu(Z)$. Note that $\partial W=\partial\nu(Z)$. By Lemma 2.1, $\partial\nu(Z)\cong\mathbb{S}(L^{\otimes k}|_{Z})=E_{Z}$. Therefore $\partial W\cong E_{Z}$. In Corollary 3, we deduced that $(W^{6},E_{Z})$ satisfies the hypothesis of Lemma 3.1 with $n=6$ and $k=3$. Therefore, $\dim_{mc}\widetilde{E_{Z}}\leq 3$.
  3. Proof of (3).
    By Proposition 5, the map $f:Z\to M$ is of positive degree. Moreover, by hypothesis $b_{2}^{+}(M)>1$. Since $f$ is of positive degree, the induced map
    $f^{*}:H^{2}(M;\mathbf{Q})\to H^{2}(Z;\mathbf{Q})$
    is an injective map of quadratic spaces, where both domain and range are equipped with the quadratic form given by the cup product $\smile$. Therefore, $b_{2}^{+}(Z)>1$. In [Tau94], Taubes proved that symplectic $4$-manifolds with $b_{2}^{+}>1$ have non-zero Seiberg-Witten invariant. Since having non-zero Seiberg-Witten invariant obstructs the existence of PSC metrics, this in particular implies $Z$ does not admit a PSC metric.
  4. Proof of (4).
    By Proposition 5, the map $f:Z\to M$ is of positive degree. Moreover, by hypothesis $M$ admits a map of positive degree to an aspherical manifold. Composing these maps, we get a positive degree map from $Z$ to an aspherical manifold. Then by [CLL23, Corollary 3], $Z$ does not admit a PSC metric.
  5. Proof of (5).
    Since $M$ is symplectically aspherical, by Proposition 6, $M$ is a rationally essential manifold. Since the map $f:Z\to M$ if of positive degree by Proposition 5, $Z$ must be rationally essential as well. Now, by Lemma 2,
    $\pi_{1}(Z)\cong\pi_{1}(M\times S^{2})\cong\pi_{1}(M).$
    By hypothesis, $\pi_{1}(M)$ is amenable. Therefore, $Z$ is a rationally essential manifold with amenable fundamental group. Hence, by [Dra11], $\dim_{mc}Z=4$.
4.2. Proof of Corollary B
Before we move to the proof of Corollary B we give several examples of $4$-manifolds $M$ satisfying the hypotheses of Theorem A.

### v3

@@ 2.3. Donaldson divisors of $M\times S_{g}$
Take any regular value $p\in M$. Then the oriented count #$\#f^{-1}(p)$ is equal to the oriented intersection number #$\#([Z]\cap[\{p\}\times S_{g}])$. Therefore, we have
$\deg(f)$ =#$=\#([Z]\cap[\{p\}\times S_{g}])$
@@
The last inequality holds because $\{p\}\times S_{g}$ is a symplectic submanifold of $(M\times S_{g},\omega)$. ∎
@@ 2.4. Symplectically aspherical manifolds
Example 6.
It is immediate from Definition 4 that symplectic submanifolds of symplectically aspherical manifolds are also symplectically aspherical. Since symplectic manifolds $(X,\omega)$ with $\pi_{2}(X)=0$ (for instance, product of surfaces of genus $g\geq 1$) are automatically symplectically aspherical, Donaldson’s theorem (Theorem 1) provides a large class of symplectically aspherical manifolds.
@@
3. Proofs of the main results
3.1. Proof of Theorem A
In this section, we shall prove our main result. We begin with a general lemma relating macroscopic dimension of a manifold with the homotopy type of a nullcobordism of the manifold.
@@ 3.1. Proof of Theorem A
To this end, let us assume without loss of generality that $f$ is a simplicial map. This can be arranged by the simplicial approximation theorem. Let $\widetilde{p}\in\widetilde{K}$ be a point lying over $p\in K$. Then, we claim $\widetilde{f}^{-1}(\widetilde{p})$ is isometric to $f^{-1}(p)$. Indeed, note that $f^{-1}(p)\hookrightarrow W$ induces the zero map on $\pi_{1}$, since otherwise there would be a loop $\gamma\subset f^{-1}(p)$ which is not nullhomotopic in $W$ such that $f_{*}[\gamma]=0\in\pi_{1}(K)$. This would be absurd, as $f$ induces an isomorphism on $\pi_{1}$. Thus, $f^{-1}(p)$ lifts homeomorphically in the cover to a unique subset in $\widetilde{Y}$ containing $\widetilde{p}$. By uniqueness, this lift is exactly $\widetilde{f}^{-1}(\widetilde{p})$. Consequently, the covering projection must be a homeomorphism:
@@
We are now ready to give a proof of Theorem A.
Proof of Theorem A.
Let $(M^{2n+2},\omega)$ be an integral symplectic manifold. Let $L$ be the pre-quantum line bundle on $(M,\omega)$. Let $Z^{2n}\subset(M,\omega)$ be a Donaldson divisor, with $[Z]=\mathrm{PD}(k\cdot[\omega])$, for $k\gg 1$ sufficiently large. We set $E_{Z}:=\mathbb{S}(L^{\otimes k}|_{Z})$ as the unit circle bundle of $L^{\otimes k}$ restricted to $Z$ (as introduced in Lemma 2.1).
  1. Proof of (1).
    Let $\nu(Z)$ be a tubular neighborhood of $Z$ in $X$. Consider the manifold with boundary $W:=M\setminus\nu(Z)$. From Lemma 2.2, $W$ has a handle decomposition consisting of handles of index $\leq n$. Let $V:=W\setminus B^{2n}$ be the $2n$-manifold obtained from removing a $2n$-ball $B^{2n}\subset W^{\circ}$ from the interior of $W$. Then $V$ is a cobordism between $S^{2n-1}$ and $\partial\nu(Z)$. But $\partial\nu(Z)=S(L^{\otimes k}|_{Z})$ by Lemma 2.1. Note that $V$ has handles of index at most $n$. Therefore, $\mathbb{S}(L^{\otimes k}|_{Z})$ is obtained from $S^{2n-1}$ by surgeries of codimension $\geq n\geq 3$. Since $S^{2n-1}$ is PSC, so is $\mathbb{S}(L^{\otimes k}|_{Z})$ by the Gromov-Lawson-Schoen-Yau surgery theorem [SY79, GL80] which states that any manifold obtained from a PSC manifold by surgeries of codimesion $\geq 3$ is also PSC.
  2. Proof of (2).
    Once again, let us consider the manifold with boundary $W:=M\setminus\nu(Z)$ where $\nu(Z)$ denotes a tubular neighborhood of $Z$ in $X$. Note that $\partial W=\partial\nu(Z)$. By Lemma 2.1, $\partial\nu(Z)\cong\mathbb{S}(L^{\otimes k}|_{Z})=E_{Z}$. Therefore $\partial W\cong E_{Z}$. In Corollary 3, we deduced that $(W^{2n+2},E_{Z})$ satisfies the hypothesis of Lemma 3.1 with $k=n+1$. Therefore, we conclude $\dim_{mc}(\widetilde{E_{Z}})\leq n+1$.
  3. Proof of (3).
    Since $(M,\omega)$ is symplectically aspherical and $Z\subset M$ is a symplectic submanifold, $Z$ is also symplectically aspherical (see, Example 6). Therefore, by Proposition 7, $Z$ is a rationally essential manifold. By Lemma 2, $\pi_{1}(Z)\cong\pi_{1}(M)$. By hypothesis, $\pi_{1}(M)$ is amenable. Therefore, $Z^{2n}$ is a rationally essential manifold with amenable fundamental group. Hence, by [Dra11, Theorem 7.2], $\dim_{mc}Z=2n$.
  4. Proof of (4).
    Here $M=N\times S_{g}$, with symplectic form $\omega=\omega_{N}\oplus\omega_{0}$, for some integral symplectic manifold $(N,\omega_{N})$ and a positive integral area form $\omega_{0}$ on $S_{g}$. Choose a Donaldson divisor $Z\subset(N\times\Sigma_{g},\omega)$ such that $[Z]=\mathrm{PD}(k\cdot[\omega])$, for some $k\gg 1$ sufficiently such that, moreover, $k$ is even. Let $f:Z\to N$ be the restriction of the projection map $M=N\times S_{g}\to N$ to $Z$. By Proposition 5 the map $f:Z\to N$ is of positive degree. By hypothesis, $N$ is enlargeable. Therefore, to show that $Z$ is also enlargeable, we just need to show that there is a finite cover $\widehat{Z}\to Z$ such that $\widehat{Z}$ is spin.
    We begin by noting that the inclusion $j:Z\hookrightarrow M$ induces an isomorphism on $\pi_{1}$, by Proposition 2. Since $N$ is enlargeable, there is a finite cover $\widehat{N}\to N$ such that $\widehat{N}$ is spin. Let $\widehat{M}:=\widehat{N}\times S_{g}$. Since $\widehat{M}$ is a product of spin manifolds, it is also spin. As $p:\widehat{M}\to M$ is a finite cover, it corresponds to some finite index subgroup $H\leq\pi_{1}(M)\cong\pi_{1}(Z)$. Let $p:\widehat{Z}\to Z$ be the finite cover of $Z$ corresponding to $H\leq\pi_{1}(Z)$. Note that the inclusion $j:Z\hookrightarrow M$ lifts canonically to an inclusion $\widehat{j}:\widehat{Z}\hookrightarrow\widehat{M}$.
    Let $\widehat{\omega}:=p^{*}\omega$ denote the lift of the symplectic form on $M$ to $\widehat{M}$. Then, $(\widehat{M},\widehat{\omega})$ is a symplectic manifold, and $\widehat{j}:\widehat{Z}\to\widehat{M}$ is a symplectic embedding. From (2.3) in Proposition 4, we have:
    $c_{1}(Z)=j^{*}c_{1}(M)-k\cdot j^{*}[\omega]$ (3.1)
    By pulling back (3.1) to the cover $p:\widehat{M}\to M$, we obtain:
    $c_{1}(\widehat{Z})=j^{*}c_{1}(\widehat{M})-k\cdot\widehat{j}^{*}[\widehat{\omega}]$ (3.2)
    Next, we reduce both sides of (3.2) modulo $2$. Recall that the first Chern class of an almost complex manifold is an integral lift of the second Stiefel-Whitney class. Since $\widehat{M}$ is spin, $w_{2}(\widehat{M})=0$. Furthermore, as $k$ is even, we conclude
    $w_{2}(\widehat{Z})\equiv c_{1}(\widehat{Z})\equiv 0\pmod{2}$
    Hence, $\widehat{Z}$ is spin. Therefore, $Z$ is enlargeable. In particular, by Gromov-Lawson [GL80], $Z$ does not admit a metric of positive scalar curvature.
Finally, it follows from Proposition 4 that for distinct, suffficiently large integers $k\neq l$, the corresponding Donaldson divisors $Z_{k},Z_{l}\subset(M,\omega)$ are not homeomorphic. Thus, one finds infinitely many manifolds $Z^{2n}$ satisfying Conclusions $(1)$-$(4)$. ∎
3.2. Proof of Theorem B
The proof of this result was communicated to us by Hanke [Han25]. We begin by stating the following proposition.
Proposition 8.
Let $M^{3}$ be a closed orientable $3$-manifold such that $M$ is either enlargeable or $\dim_{mc}(\widetilde{M})=3$. Then the Kneser-Milnor prime decomposition of $M$ must contain an aspherical summand.
Proof.
We prove the contrapositive. Let $M^{3}$ be a closed $3$-manifold. Let
$M=P_{1}\#P_{2}\#\cdots\#P_{n}$
be the prime decomposition of $M$, and suppose $P_{i}$ is not aspherical for any $1\leq i\leq n$. It follows from the sphere theorem that every closed orientable irreducible $3$-manifold with infinite fundamental group is aspherical. Therefore, either $P_{i}$ is prime but not irreducible so that $P_{i}\cong S^{1}\times S^{2}$, or $P_{i}$ has finite fundamental group. Consequently, we must have $\pi_{1}(M)\cong F_{k}*G_{1}*\cdots*G_{\ell}$ where $G_{i}$ is finite for all $1\leq i\leq\ell$ and $F_{k}$ is the free group of rank $k$. Therefore,
$K(\pi_{1}(M),1)\cong\bigvee_{i=1}^{k}S^{1}\vee\bigvee_{i=1}^{\ell}K(G_{i},1)$
Since $G_{i}$ is a finite group, $\widetilde{H}_{*}(K(G_{i},1);\mathbf{Q})=0$. Thus, $H_{3}(K(\pi_{1}(M),1);\mathbf{Q})=0$. Therefore, $M$ can not be rationally essential. By [HS06, Theorem 1.4], closed enlargeable manifolds are rationally essential. Therefore, $M$ can not be enlargeable either.
Next, we show $\dim_{mc}(\widetilde{M})\neq 3$. Consider the homomorphism
$\varphi:\pi_{1}(M)\cong F_{k}*G_{1}*\cdots*G_{\ell}\to G_{1}\times\cdots\times G_{\ell},$
given by quotienting out the free factor $F_{k}$, followed by abelianization. Since $\varphi$ is surjective and $G_{1}\times\cdots\times G_{\ell}$ is finite, $\ker(\varphi)$ must be a finite index subgroup of $\pi_{1}(M)$. It follows from the Kurosh subgroup theorem that $\ker(\varphi)$ is free of finite rank. Let $\widehat{M}\to M$ be the finite sheeted cover corresponding to $\ker(\varphi)$.
Since $\pi_{1}(\widehat{M})\cong\ker(\varphi)$ is free, any prime summand of $\widehat{M}$ is a prime $3$-manifold with free fundamental group. Note that there are no closed orientable irreducible $3$-manifold with free fundamental group. Indeed, such a $3$-manifold must be aspherical by the sphere theorem. But this implies the cohomological dimension of a free group is $3$, which is absurd. Therefore, all nontrivial prime summands of $\widehat{M}$ must be prime but not irreducible $3$-manifolds, i.e. $S^{1}\times S^{2}$. Therefore, $\widehat{M}\cong\#^{m}(S^{1}\times S^{2})$.
Next, observe that the universal cover of $M$ agrees with the universal cover of $\widehat{M}$. However, the universal cover of $\widehat{M}\cong\#^{m}(S^{1}\times S^{2})$ is obtained from the standard Cayley graph of $F_{m}$ by assigning vertex spaces as $S^{3}$, edge spaces as $S^{2}\times I$, and tubing the two vertex spaces corresponding to the endpoints of any given edge by an $S^{2}\times I$. Consider the map $f:\widetilde{M}\to Cay(F_{k})$ obtained by collapsing down the vertex spaces to the vertices and edge spaces to the edges of the Cayley graph. Thus, the fibers of $f$ have uniformly bounded diameter. Therefore, $\dim_{mc}(\widetilde{M})\leq 1$. In particular, $\dim_{mc}(\widetilde{M})\neq 3$, as desired. ∎
We now proceed to the proof of Theorem B:
Proof of Theorem B.
Let $M^{3}$ be a closed $3$-manifold such that $M$ is either enlargeable or $\dim_{mc}(\widetilde{M})=3$, and let $E\to M$ be a circle bundle over $M$ such that the total space of $E$ admits a PSC metric. As a first step, we argue that it can be ensured without loss of generality that $(1)$ $M$ is orientable, $(2)$ the circle bundle $E\to M$ is fiberwise orientable.
@@ 3.2. Proof of Theorem B
    If $M$ is enlargeable, then $M$ is orientable by definition. If $\dim_{mc}(\widetilde{M})=3$ and $M$ is not orientable, we pass to the orientation double cover of $M$ and pull the bundle $E\to M$ back to this cover. As the macroscopic dimension of a manifold is a quantity that only depends on the universal cover of the manifold, the cover continues to have macroscopic dimension $3$. Moreover, the total space of the pullback of $E\to M$ is a double cover of the total space of $E$. Therefore, the total space of the pulled back bundle also admits a PSC metric. Therefore, without loss of generality, we may assume $M$ is orientable.
@@
    If the circle bundle $E$ is not fiberwise orientable, then $w_{1}(E)\neq 0\in H^{1}(M;\mathbf{Z}_{2})$. By using the isomorphism $H^{1}(M;\mathbf{Z}_{2})\cong\mathrm{Hom}(\pi_{1}(M),\mathbf{Z}_{2}),$ we obtain a homomorphism $\phi:\pi_{1}(M)\to\mathbf{Z}_{2}$ corresponding to $w_{1}(E)$. We pass to the double cover of $M$ corresponding to the index $2$ subgroup $\ker(\phi)$ of $\pi_{1}(M)$. The pullback of $E$ to this cover is an oriented circle bundle by construction. By the reasoning in Part $(1)$, the total space of the pulled back bundle continues to be PSC. If $\dim_{mc}(\widetilde{M})=3$, then the the macroscopic dimension of this cover is also $3$. If $M$ is enlargeable, then the cover is also enlargeable: it admits a degree $2$ map to $M$ and it is spin since all orientable $3$-manifolds are spin. Hence, the cover is also enlargeable. Therefore, we may also assume without loss of generality that the bundle $E\to M$ is fiberwise orientable.
By Proposition 8, $M$ contains an aspherical prime summand $K$, so that $M=K\#N$ for some $3$-manifold $N$. Recall that fiberwise oriented circle bundles $E\to M$ are classified by the Euler class $e(E)\in H^{2}(M;\mathbf{Z})$. Note that
$H^{2}(M;\mathbf{Z})\cong H^{2}(K;\mathbf{Z})\oplus H^{2}(N;\mathbf{Z})$
Therefore, $E$ is isomorphic to a fiber connect sum of two circle bundles over $K$ and $N$, respectively. Let $L\to K$ denote the circle bundle corresponding to the aspherical summand $K$. Therefore, the total space of $E$ admits a degree $1$ map to the total space of $L$. Since the total space of a circle bundle over an aspherical manifold is aspherical, this implies that the total space of $E$ admits a degree $1$ map to an aspherical $4$-manifold. Hence, the total space of $E$ can not admit a positive scalar curvature metric by a result of Chodosh, Li and Liokumovich [CLL23, Corollary 3]. This concludes the proof. ∎
4. Further examples and remarks
In this section, we provide several examples of manifolds satisfying the conclusions of Theorem A. We begin with the simplest example given by a symplectic torus:
Example 9 (Torus).
Let $M=T^{2n}$ for $n\geq 3$ and let $\omega$ be the standard symplectic form on $T^{2n}=\mathbf{R}^{2n}/\mathbf{Z}^{2n}$, normalized to be integral. Since $\pi_{2}(T^{2n})=0$, $(M,\omega)$ is a symplectically aspherical manifold. Furthermore, $\pi_{1}(T^{2n})=\mathbf{Z}^{2n}$ is an amenable group. Moreover, $M=T^{2n-2}\times T^{2}$ as symplectic manifolds, and $T^{2n-2}$ is enlargeable. Therefore, $(M,\omega)$ satisfies the hypothesis of Parts $(3)$ and $(4)$ of Theorem A.
We remark that using Example 9, we can produce examples of macroscopically large manifolds $Z$ admitting circle bundles $p:E\to Z$ such that the total space $E$ is macroscopically small, yet the asymptotic dimension of $\pi_{1}(E)$ is large. This indicates that the macroscopic smallness of $E$ cannot be deduced from just the fundamental group. [^4] For any metric space $X$, $\dim_{mc}(X)\leq\mathrm{asdim}(X)$. For more information on the asymptotic dimension, we refer the reader to the survey [BD08].